# Multimodal Representation Learning and Fusion thanks:

### v1

@@ V Applications of Multi-modal
The definition of multi-modality has gone through tremendous change over the past few years, with novel definitions emerging to suit the era of machine learning [Parcalabescu et al., 2021]. Multi-modal learning has seen great success in applications like computer vision, natural language processing, and speech recognition [Cohn et al., 2024]. Recent developments in Multi-modal Large Language Models (MLLMs) have brought improvements to tasks like vision-language pre-training and representation learning [Gan et al., 2022]. New types of data fusion, including early and late fusion, have enhanced efficiency [Cohn et al., 2024], as formalized in (4) and (5). However, challenges remain, such as the need for better fusion techniques and large-scale datasets [Pattnayak et al., 2024].
Recent developments in MLLMs have brought dramatic improvements to numerous tasks, including vision-language pre-training [Gan et al., 2022] and multi-modal representation learning [Manzoor et al., 2023]. The application of transformer-based models has been particularly successful at reaching state-of-the-art levels in numerous multi-modal tasks [Wang et al., 2024]. Still, there are challenges and limitations despite recent advances in multi-modal learning. For instance, the necessity for better fusion techniques to amalgamate data from various modalities is one of the major challenges [Cohn et al., 2024]. Furthermore, the creation of large-scale data sets that can be used to train and test multi-modal models is still an existing challenge [Pattnayak et al., 2024].
Recent years have seen major advances in both the methodologies and the technical details involved in multi-modal learning. New types of data fusion, including mid-fusion and late-fusion, have enhanced the efficiency of multi-modal models [Cohn et al., 2024]. Graph-based methods have similarly been applied to strengthen reviews of literature as well as to enhance the analysis of multi-modal data [Cohn et al., 2024]. In addition, pretraining objectives, including masked language models and next sentence prediction, have been found to be useful to enhance the performance of MLLMs [Wang et al., 2024].
And yet, even with these gains, the future of multi-modal learning remains full of questions, and opportunities. A lot of attention is now focused on designing smarter and more adaptable model architectures that can handle a wide range of modalities smoothly and efficiently [Manzoor et al., 2023]. In addition, improving our understanding of how different modalities interact with each other, especially in real-time or noisy settings, remains a significant challenge [Cohn et al., 2024]. It is not only about merging signals; it is about understanding their relationships and using that understanding to make better decisions.
@@
Looking forward, we can expect to see even more multi-modal applications in traditional areas like natural language processing, computer vision, and speech recognition. These areas are already mature, but multi-modality brings new possibilities. Ultimately, the long-term goal of multi-modal learning is to create systems that can perform at a level similar to humans on complex, real-world tasks. Systems can see, listen, read, and understand all at once [Pattnayak et al., 2024]. That goal may still be far away, but with each new model, dataset, or training method, we get a little bit closer.
@@ VI Multi-modal Neural Architecture Search
Multi-modal neural architecture search (NAS) automates the design of efficient neural networks for multi-modal tasks [Yu et al., 2020]. Methods like gradient-based NAS and sequential model-based exploration have achieved state-of-the-art results [Pérez-Rúa et al., 2019]. However, challenges such as computational cost and generalizability persist [Wu et al., 2024]. Fusion strategies, including early and deep fusion, are critical to NAS performance [Akbari et al., 2023].
@@
One predominant theme that is found throughout the literature is the automation of architectural design, aiming to save manual efforts as well as to increase performance [Yu et al., 2020][Pérez-Rúa et al., 2019]. Most methods try to offer flexible frameworks that can be capitalized upon across various multi-modal tasks as well as data sets [Yin et al., 2022][Wadekar et al., 2024]. The efficiency of the search process is another target for several researchers to minimize the computational cost as well as the time consumed by NAS [Pasunuru and Bansal, 2019][Alletto et al., 2020].
@@
Although these progresses have been made, there exist differing opinions and methods within this domain. To be specific, [Pérez-Rúa et al., 2019] is interested in this novel search space exclusively for fusion architectures, whereas others, such as [Yu et al., 2020] and [Yin et al., 2022] investigate more comprehensive frameworks that support multiple different operations and tasks. The modalities’ integration is even addressed differently between studies, where some prioritize early fusion [Akkus et al., 2023], while others propose deep fusion or the fusion of both [Akbari et al., 2023]. In addition, the application of pretrained models differs, with some methods utilizing pretrained unimodal backbones [Yin et al., 2022], where other methods rely on training from scratch or alternative pretraining methods such as [Akbari et al., 2023]’s Masked Architecture Modeling (MAM).
@@
The technical methodologies employed in multi-modal NAS are diverse and include gradient-based NAS [Yu et al., 2020], sequential model-based exploration [Pérez-Rúa et al., 2019], bilevel searching schemes [Yin et al., 2022], and bi-modal learning [Akbari et al., 2023]. These approaches have been applied to various multi-modal tasks, including those involving vision, language, and audio modalities [Baltrušaitis et al., 2018][Manzoor et al., 2023]. However, despite these advancements, challenges persist, including the computational cost associated with NAS [Wu et al., 2024], ensuring generalizability across tasks [Liang et al., 2024c], interpreting and explaining the performance of searched architectures [McKinzie et al., 2023], and scalability to new modalities [Wei and Hu, 2024].
@@
In summary, multi-modal neural architecture search has emerged as a dynamic and fast-developing field with great potential to improve both the efficiency and effectiveness of multi-modal learning systems. While a number of recent contributions have made meaningful steps toward addressing issues such as cost, transferability, and architectural complexity, much remains to be explored. Questions related to scalability, interpretability, and generalization are still open and continue to limit the full realization of NAS in practical multi-modal applications [Yin et al., 2022][Wadekar et al., 2024]. Looking forward, it is reasonable to expect that more advanced frameworks and algorithms will be introduced to meet these challenges. With further work, we may see models that are not only better suited to handling the unique difficulties of multi-modal learning, but also more capable of delivering consistent performance across a wide array of real-world tasks [Akbari et al., 2023][Wadekar et al., 2024].
@@ VII Real-Time and Low-Resource Scenarios
Efficient multi-modal processing has long been a central concern in multi-modal learning, especially under real-time and low-resource conditions. Researchers have focused on integrating information from different modalities in ways that are both accurate and highly efficient. For instance, the survey paper “Multi-modal Co-learning: Challenges, Applications with Datasets, Recent Advances and Future Directions” [Rahate et al., 2022] offers a comprehensive review of multi-modal co-learning, highlighting issues like noisy or missing modalities and proposing a taxonomy to structure the problem. Robustness is essential in such scenarios, and attribution regularization encourages models to utilize all available modalities, mitigating single-modality dominance [Yerramilli et al., 2024].
@@
where $\mathcal{L}_{\text{task}}$ is the task-specific loss, $\frac{\partial\mathcal{L}_{\text{task}}}{\partial\mathbf{x}_{m}}$ measures modality $m$’s gradient contribution, and $\lambda$ controls regularization strength. Additionally, greedy modality selection improves efficiency under computational limits by prioritizing informative modalities [Cheng et al., 2022]. These approaches reflect the field’s emphasis on clear definitions, structured analysis, and practical challenges.
Another important contribution comes from the paper “Greedy Modality Selection via Approximate Submodular Maximization” [Cheng et al., 2022]. It introduces a state-of-the-art approach with an efficient algorithm designed to select the most informative and complementary modalities under computational limits. The method shows strong results not only on synthetic tasks but also in real-world datasets, making a clear case for modality optimization as a practical way to improve processing efficiency. By contrast, the work titled “Modular and Parameter-Efficient Multi-modal Fusion with Prompting” [Liang et al., 2022] proposes a lightweight solution using modal prompts and novel vector alignment. The technique performs comparably to other mainstream methods even under low-resource scenarios. This shows that prompting, though relatively new in this field, can be a flexible and scalable approach for multi-modal learning when resources are limited.
Robustness is another essential topic that has gained attention in recent work. One representative study, “On Robustness in Multi-modal Learning” [McKinzie et al., 2023], proposes a framework to identify the weaknesses in commonly used representation learning techniques. It further suggests targeted interventions to improve robustness. In a related direction, the paper “Attribution Regularization for Multi-modal Paradigms” [Yerramilli et al., 2024] proposes a new regularization term. This term encourages models to better utilize all available modalities, which also helps mitigate the problem of single-modality dominance. These contributions suggest that improving generalizability and robustness is not only possible but necessary, especially as models are applied to real-world environments where data sources may be uncertain or unstable.
A common thread running through these contributions is the growing recognition of efficiency and scalability as core needs in multi-modal learning. For example, the work “Multi-modal Information Bottleneck: Learning Minimalist Models” [Mai et al., 2022] emphasizes the value of minimalist modeling, how to reduce redundancy while still maintaining performance. This also connects to broader questions around evaluation: what counts as “good enough” performance in constrained settings, and how should it be measured across different tasks and systems?
@@ VIII AutoML for Multi-modal Learning: Recent Developments and Future Directions
The integration of automated machine learning (AutoML) with multi-modal learning has advanced rapidly, driven by increasing data complexity and the demand for efficient, user-friendly solutions [Luo et al., 2024, Tang et al., 2024a]. Frameworks like AutoM3L, leveraging large language models, automate multi-modal training pipelines, reducing manual effort [Luo et al., 2024], while AutoGluon-multi-modal (AutoMM) enables fine-tuning of foundation models with minimal code, supporting diverse tasks [Tang et al., 2024a]. Simple techniques, such as cross-modal alignment, yield robust results [Tang et al., 2024b], as shown in (2). However, challenges like pipeline complexity and inconsistent evaluation metrics persist [Shi et al., 2021].
@@
The combination of AutoML with computer vision has also been widely studied. Because data is getting bigger and more complex, researchers try to make AutoML easier to use. AutoM3L [Luo et al., 2024] provides an LLM-based solution to help users set up multi-modal pipelines without much effort. AutoMM [Tang et al., 2024a] also gives a simple and efficient way to handle different modalities and improve performance on many real-world tasks.
@@
Recent studies tell us something very clear: handling multi-modal data well is no longer a bonus, it is a must. The paper “Bag of Tricks for Multi-modal AutoML” [Tang et al., 2024b] runs a large benchmark that covers images, texts, and tables. It shows that even simple tricks, such as fusion, data augmentation, and cross-modal alignment, can bring strong and stable results. But this paper also reminds us that mixing multiple data types is not an easy job. So we need AutoML tools that can link different modalities together quickly and smartly.
@@
Another big trend is the use of advanced models, especially large language models (LLMs) and foundation models [Luo et al., 2024][Tang et al., 2024a]. These models make it possible to automate very complicated tasks, improve overall results, and let more people use machine learning, even if they are not experts. The shift from rule-based tools to model-based methods shows that the community now trusts LLMs to do reasoning, generate code, and provide useful interactions.
@@
However, even with this exciting progress, there are still many problems. First, since we are dealing with many types of data, the pipelines become heavier and slower. Also, it is still hard to make the models explainable [Tuggener et al., 2019][Shen et al., 2018]. Another issue is evaluation. Right now, different studies use different metrics and datasets. So we cannot compare fairly. One common benchmark is badly needed to help build strong and general AutoML systems [Shi et al., 2021].
@@
Looking back at how AutoML has changed, we can see a big shift. The paper “Automated Machine Learning , A Brief Review” [Escalante, 2021] explains this clearly. In the early years, AutoML focused on tuning hyperparameters or picking models. But today, researchers care more about how to use AutoML in real life, especially with messy and mixed multi-modal data. People want tools that are not only fast but also easy to use and able to handle real-world problems.
@@
To sum up, when AutoML meets multi-modal learning, things get very interesting. We now have many strong tools and smart frameworks, but we also face a lot of challenges behind the scenes. If we want to unlock the full power of AutoML, we still need to solve key pain points, like pipeline complexity, scaling up, explainability, and fair benchmarking. Fixing these will make ML tools faster, more friendly, and strong enough to handle difficult tasks. Also, this research is closely linked to the big direction of AI, since more and more LLMs and foundation models are being developed [Luo et al., 2024][Tang et al., 2024a]. So we believe continuous research and teamwork are the key to pushing AutoML forward in the multi-modal world.
@@ IX Unsupervised and Semi-Supervised Multi-modal Learning
Unsupervised and semi-supervised learning are among the fastest-growing directions in multi-modal research, driven by their ability to handle unseen modality combinations at inference time [Zhang et al., 2023]. Recent work emphasizes task-specific modality selection, prioritizing utility over comprehensive modeling [Parcalabescu et al., 2021]. Theoretical bounds further quantify cross-modal interactions in semi-supervised scenarios [Liang et al., 2024b].
@@
where $I(\mathbf{X}_{1};\mathbf{X}_{2}\mid\mathbf{Y})$ is the conditional mutual information, $H(\cdot)$ denotes entropy, and $H(\mathbf{Y}\mid\mathbf{X}_{1},\mathbf{X}_{2})$ is the conditional entropy. Teacher-student frameworks also enhance generalization without labeled data [Xue et al., 2021].
@@
At the same time, there is an ongoing effort to understand whether different modalities are genuinely interacting, or just existing side by side. Liang et al. [Liang et al., 2024b] proposed a set of theoretical bounds to estimate the interaction level in semi-supervised scenarios. Sure, these bounds do not offer precise yes-or-no answers, but they do help guide us in both evaluating and designing better systems. Meanwhile, another interesting work [Xue et al., 2021] introduced a teacher-student framework where a pre-trained single-modal teacher can still pass meaningful knowledge to a multi-modal student, even in the absence of labeled data. Surprisingly, this approach improves generalization and lowers the model’s sensitivity to noisy signals.
@@
If there is one key message across these studies, it is this: flexibility and robustness are no longer optional, they are essential. Today’s models need to be ready for missing data, noisy inputs, or strange, previously unseen modality mixes. What makes this possible? Strong representation learning. Many teams are now using self-supervised or unlabeled data to learn meaningful features; two good examples can be found in [Akkus et al., 2023][Cohn et al., 2024]. These works push the field forward not just in terms of technical depth, but also in how we understand multi-modal systems in real-world or even complex environmental contexts.
@@
However, not everyone agrees on what “multi-modality” should mean. Some papers, like [Parcalabescu et al., 2021], advocate for task-specific definitions, meaning the way we define and handle modalities depends on what the task actually needs. Others hold on to more traditional, general-purpose views. Even worse, there is no unified standard for measuring cross-modal interactions. One paper turns to information theory [Liang et al., 2024b]; others prefer to rely on practical benchmarks or novel metrics. So depending on the lens you use, the evaluation result may vary dramatically.
@@
Technically, the field is vibrant with innovation. For example, [Zhang et al., 2023] explores novel mixes of projection and fusion. Liang et al. [Liang et al., 2024b] give mathematical tools, bounds and ranges, to quantify how well different modalities actually interact. [Xue et al., 2021] makes the case that a model can still learn without labels, using cross-modal knowledge transfer. On top of that, others like [Chen et al., 2021][Gao et al., 2025] are testing semi-supervised setups in tasks like 3D object recognition, or working on ways to reduce biases introduced by imbalanced modality signals.
@@
Of course, many problems are still far from solved. One major pain point is data, large labeled datasets remain costly and time-consuming to collect [Baltrušaitis et al., 2018][Rahate et al., 2022]. Another issue is, we still do not have a reliable way to estimate cross-modal interaction without supervision [Liang et al., 2024b]. And let’s be honest: we are still waiting for that one model that can truly handle any level of supervision and still be robust in the face of messy, real-world input.

### v2

@@ V Applications of Multi-modal
The definition of multi-modality has gone through tremendous change over the past few years, with novel definitions emerging to suit the era of machine learning [2103.06304]. Multi-modal learning has seen great success in applications like computer vision, natural language processing, and speech recognition [2408.14491]. Recent developments in Multi-modal Large Language Models (MLLMs) have brought improvements to tasks like vision-language pre-training and representation learning [2210.09263]. New types of data fusion, including early and late fusion, have enhanced efficiency [2408.14491], as formalized in (4) and (5). However, challenges remain, such as the need for better fusion techniques and large-scale datasets [2412.17759].
Recent developments in MLLMs have brought dramatic improvements to numerous tasks, including vision-language pre-training [2210.09263] and multi-modal representation learning [2302.00389]. The application of transformer-based models has been particularly successful at reaching state-of-the-art levels in numerous multi-modal tasks [2408.01319]. Still, there are challenges and limitations despite recent advances in multi-modal learning. For instance, the necessity for better fusion techniques to amalgamate data from various modalities is one of the major challenges [2408.14491]. Furthermore, the creation of large-scale data sets that can be used to train and test multi-modal models is still an existing challenge [2412.17759].
Recent years have seen major advances in both the methodologies and the technical details involved in multi-modal learning. New types of data fusion, including mid-fusion and late-fusion, have enhanced the efficiency of multi-modal models [2408.14491]. Graph-based methods have similarly been applied to strengthen reviews of literature as well as to enhance the analysis of multi-modal data [2408.14491]. In addition, pretraining objectives, including masked language models and next sentence prediction, have been found to be useful to enhance the performance of MLLMs [2408.01319].
And yet, even with these gains, the future of multi-modal learning remains full of questions, and opportunities. A lot of attention is now focused on designing smarter and more adaptable model architectures that can handle a wide range of modalities smoothly and efficiently [2302.00389]. In addition, improving our understanding of how different modalities interact with each other, especially in real-time or noisy settings, remains a significant challenge [2408.14491]. It is not only about merging signals; it is about understanding their relationships and using that understanding to make better decisions.
@@
Looking forward, we can expect to see even more multi-modal applications in traditional areas like natural language processing, computer vision, and speech recognition. These areas are already mature, but multi-modality brings new possibilities. Ultimately, the long-term goal of multi-modal learning is to create systems that can perform at a level similar to humans on complex, real-world tasks. Systems can see, listen, read, and understand all at once [2412.17759]. That goal may still be far away, but with each new model, dataset, or training method, we get a little bit closer.
@@ VI Multi-modal Neural Architecture Search
Multi-modal neural architecture search (NAS) automates the design of efficient neural networks for multi-modal tasks [2004.12070]. Methods like gradient-based NAS and sequential model-based exploration have achieved state-of-the-art results [1903.06496]. However, challenges such as computational cost and generalizability persist [2409.07825]. Fusion strategies, including early and deep fusion, are critical to NAS performance [2310.17737].
@@
One predominant theme that is found throughout the literature is the automation of architectural design, aiming to save manual efforts as well as to increase performance [2004.12070][1903.06496]. Most methods try to offer flexible frameworks that can be capitalized upon across various multi-modal tasks as well as data sets [2104.09379][2405.17927]. The efficiency of the search process is another target for several researchers to minimize the computational cost as well as the time consumed by NAS [1906.05226][2003.01181].
@@
Although these progresses have been made, there exist differing opinions and methods within this domain. To be specific, [1903.06496] is interested in this novel search space exclusively for fusion architectures, whereas others, such as [2004.12070] and [2104.09379] investigate more comprehensive frameworks that support multiple different operations and tasks. The modalities’ integration is even addressed differently between studies, where some prioritize early fusion [2301.04856], while others propose deep fusion or the fusion of both [2310.17737]. In addition, the application of pretrained models differs, with some methods utilizing pretrained unimodal backbones [2104.09379], where other methods rely on training from scratch or alternative pretraining methods such as [2310.17737]’s Masked Architecture Modeling (MAM).
@@
The technical methodologies employed in multi-modal NAS are diverse and include gradient-based NAS [2004.12070], sequential model-based exploration [1903.06496], bilevel searching schemes [2104.09379], and bi-modal learning [2310.17737]. These approaches have been applied to various multi-modal tasks, including those involving vision, language, and audio modalities [1705.09406][2302.00389]. However, despite these advancements, challenges persist, including the computational cost associated with NAS [2409.07825], ensuring generalizability across tasks [2209.03430], interpreting and explaining the performance of searched architectures [2304.04385], and scalability to new modalities [2405.17730].
@@
In summary, multi-modal neural architecture search has emerged as a dynamic and fast-developing field with great potential to improve both the efficiency and effectiveness of multi-modal learning systems. While a number of recent contributions have made meaningful steps toward addressing issues such as cost, transferability, and architectural complexity, much remains to be explored. Questions related to scalability, interpretability, and generalization are still open and continue to limit the full realization of NAS in practical multi-modal applications [2104.09379][2405.17927]. Looking forward, it is reasonable to expect that more advanced frameworks and algorithms will be introduced to meet these challenges. With further work, we may see models that are not only better suited to handling the unique difficulties of multi-modal learning, but also more capable of delivering consistent performance across a wide array of real-world tasks [2310.17737][2405.17927].
@@ VII Real-Time and Low-Resource Scenarios
Efficient multi-modal processing has long been a central concern in multi-modal learning, especially under real-time and low-resource conditions. Researchers have focused on integrating information from different modalities in ways that are both accurate and highly efficient. For instance, the survey paper “Multi-modal Co-learning: Challenges, Applications with Datasets, Recent Advances and Future Directions” [2107.13782] offers a comprehensive review of multi-modal co-learning, highlighting issues like noisy or missing modalities and proposing a taxonomy to structure the problem. Robustness is essential in such scenarios, and attribution regularization encourages models to utilize all available modalities, mitigating single-modality dominance [2404.02359].
@@
where $\mathcal{L}_{\text{task}}$ is the task-specific loss, $\frac{\partial\mathcal{L}_{\text{task}}}{\partial\mathbf{x}_{m}}$ measures modality $m$’s gradient contribution, and $\lambda$ controls regularization strength. Additionally, greedy modality selection improves efficiency under computational limits by prioritizing informative modalities [2210.12562]. These approaches reflect the field’s emphasis on clear definitions, structured analysis, and practical challenges.
Another important contribution comes from the paper “Greedy Modality Selection via Approximate Submodular Maximization” [2210.12562]. It introduces a state-of-the-art approach with an efficient algorithm designed to select the most informative and complementary modalities under computational limits. The method shows strong results not only on synthetic tasks but also in real-world datasets, making a clear case for modality optimization as a practical way to improve processing efficiency. By contrast, the work titled “Modular and Parameter-Efficient Multi-modal Fusion with Prompting” [2203.08055] proposes a lightweight solution using modal prompts and novel vector alignment. The technique performs comparably to other mainstream methods even under low-resource scenarios. This shows that prompting, though relatively new in this field, can be a flexible and scalable approach for multi-modal learning when resources are limited.
Robustness is another essential topic that has gained attention in recent work. One representative study, “On Robustness in Multi-modal Learning” [2304.04385], proposes a framework to identify the weaknesses in commonly used representation learning techniques. It further suggests targeted interventions to improve robustness. In a related direction, the paper “Attribution Regularization for Multi-modal Paradigms” [2404.02359] proposes a new regularization term. This term encourages models to better utilize all available modalities, which also helps mitigate the problem of single-modality dominance. These contributions suggest that improving generalizability and robustness is not only possible but necessary, especially as models are applied to real-world environments where data sources may be uncertain or unstable.
A common thread running through these contributions is the growing recognition of efficiency and scalability as core needs in multi-modal learning. For example, the work “Multi-modal Information Bottleneck: Learning Minimalist Models” [mai2022multimodal] emphasizes the value of minimalist modeling, how to reduce redundancy while still maintaining performance. This also connects to broader questions around evaluation: what counts as “good enough” performance in constrained settings, and how should it be measured across different tasks and systems?
@@ VIII AutoML for Multi-modal Learning: Recent Developments and Future Directions
The integration of automated machine learning (AutoML) with multi-modal learning has advanced rapidly, driven by increasing data complexity and the demand for efficient, user-friendly solutions [2408.00665, 2404.16233]. Frameworks like AutoM3L, leveraging large language models, automate multi-modal training pipelines, reducing manual effort [2408.00665], while AutoGluon-multi-modal (AutoMM) enables fine-tuning of foundation models with minimal code, supporting diverse tasks [2404.16233]. Simple techniques, such as cross-modal alignment, yield robust results [2412.16243], as shown in (2). However, challenges like pipeline complexity and inconsistent evaluation metrics persist [2111.02705].
@@
The combination of AutoML with computer vision has also been widely studied. Because data is getting bigger and more complex, researchers try to make AutoML easier to use. AutoM3L [2408.00665] provides an LLM-based solution to help users set up multi-modal pipelines without much effort. AutoMM [2404.16233] also gives a simple and efficient way to handle different modalities and improve performance on many real-world tasks.
@@
Recent studies tell us something very clear: handling multi-modal data well is no longer a bonus, it is a must. The paper “Bag of Tricks for Multi-modal AutoML” [2412.16243] runs a large benchmark that covers images, texts, and tables. It shows that even simple tricks, such as fusion, data augmentation, and cross-modal alignment, can bring strong and stable results. But this paper also reminds us that mixing multiple data types is not an easy job. So we need AutoML tools that can link different modalities together quickly and smartly.
@@
Another big trend is the use of advanced models, especially large language models (LLMs) and foundation models [2408.00665][2404.16233]. These models make it possible to automate very complicated tasks, improve overall results, and let more people use machine learning, even if they are not experts. The shift from rule-based tools to model-based methods shows that the community now trusts LLMs to do reasoning, generate code, and provide useful interactions.
@@
However, even with this exciting progress, there are still many problems. First, since we are dealing with many types of data, the pipelines become heavier and slower. Also, it is still hard to make the models explainable [1907.08392][1810.13306]. Another issue is evaluation. Right now, different studies use different metrics and datasets. So we cannot compare fairly. One common benchmark is badly needed to help build strong and general AutoML systems [2111.02705].
@@
Looking back at how AutoML has changed, we can see a big shift. The paper “Automated Machine Learning , A Brief Review” [2008.08516] explains this clearly. In the early years, AutoML focused on tuning hyperparameters or picking models. But today, researchers care more about how to use AutoML in real life, especially with messy and mixed multi-modal data. People want tools that are not only fast but also easy to use and able to handle real-world problems.
@@
To sum up, when AutoML meets multi-modal learning, things get very interesting. We now have many strong tools and smart frameworks, but we also face a lot of challenges behind the scenes. If we want to unlock the full power of AutoML, we still need to solve key pain points, like pipeline complexity, scaling up, explainability, and fair benchmarking. Fixing these will make ML tools faster, more friendly, and strong enough to handle difficult tasks. Also, this research is closely linked to the big direction of AI, since more and more LLMs and foundation models are being developed [2408.00665][2404.16233]. So we believe continuous research and teamwork are the key to pushing AutoML forward in the multi-modal world.
@@ IX Unsupervised and Semi-Supervised Multi-modal Learning
Unsupervised and semi-supervised learning are among the fastest-growing directions in multi-modal research, driven by their ability to handle unseen modality combinations at inference time [2306.12795]. Recent work emphasizes task-specific modality selection, prioritizing utility over comprehensive modeling [2103.06304]. Theoretical bounds further quantify cross-modal interactions in semi-supervised scenarios [2306.04539].
@@
where $I(\mathbf{X}_{1};\mathbf{X}_{2}\mid\mathbf{Y})$ is the conditional mutual information, $H(\cdot)$ denotes entropy, and $H(\mathbf{Y}\mid\mathbf{X}_{1},\mathbf{X}_{2})$ is the conditional entropy. Teacher-student frameworks also enhance generalization without labeled data [2103.14431].
@@
At the same time, there is an ongoing effort to understand whether different modalities are genuinely interacting, or just existing side by side. Liang et al. [2306.04539] proposed a set of theoretical bounds to estimate the interaction level in semi-supervised scenarios. Sure, these bounds do not offer precise yes-or-no answers, but they do help guide us in both evaluating and designing better systems. Meanwhile, another interesting work [2103.14431] introduced a teacher-student framework where a pre-trained single-modal teacher can still pass meaningful knowledge to a multi-modal student, even in the absence of labeled data. Surprisingly, this approach improves generalization and lowers the model’s sensitivity to noisy signals.
@@
If there is one key message across these studies, it is this: flexibility and robustness are no longer optional, they are essential. Today’s models need to be ready for missing data, noisy inputs, or strange, previously unseen modality mixes. What makes this possible? Strong representation learning. Many teams are now using self-supervised or unlabeled data to learn meaningful features; two good examples can be found in [2301.04856][2408.14491]. These works push the field forward not just in terms of technical depth, but also in how we understand multi-modal systems in real-world or even complex environmental contexts.
@@
However, not everyone agrees on what “multi-modality” should mean. Some papers, like [2103.06304], advocate for task-specific definitions, meaning the way we define and handle modalities depends on what the task actually needs. Others hold on to more traditional, general-purpose views. Even worse, there is no unified standard for measuring cross-modal interactions. One paper turns to information theory [2306.04539]; others prefer to rely on practical benchmarks or novel metrics. So depending on the lens you use, the evaluation result may vary dramatically.
@@
Technically, the field is vibrant with innovation. For example, [2306.12795] explores novel mixes of projection and fusion. Liang et al. [2306.04539] give mathematical tools, bounds and ranges, to quantify how well different modalities actually interact. [2103.14431] makes the case that a model can still learn without labels, using cross-modal knowledge transfer. On top of that, others like [2110.11601][2501.01240] are testing semi-supervised setups in tasks like 3D object recognition, or working on ways to reduce biases introduced by imbalanced modality signals.
@@
Of course, many problems are still far from solved. One major pain point is data, large labeled datasets remain costly and time-consuming to collect [1705.09406][2107.13782]. Another issue is, we still do not have a reliable way to estimate cross-modal interaction without supervision [2306.04539]. And let’s be honest: we are still waiting for that one model that can truly handle any level of supervision and still be robust in the face of messy, real-world input.